# PoIFusion: Multi-Modal 3D Object Detection via Fusion at Points of Interest

### v1

@@ Abstract
In this work, we present PoIFusion, a simple yet effective multi-modal 3D object detection framework to fuse the information of RGB images and LiDAR point clouds at the point of interest (abbreviated as PoI). Technically, our PoIFusion follows the paradigm of query-based object detection, formulating object queries as dynamic 3D boxes. The PoIs are adaptively generated from each query box on the fly, serving as the keypoints to represent a 3D object and play the role of basic units in multi-modal fusion. Specifically, we project PoIs into the view of each modality to sample the corresponding feature and integrate the multi-modal features at each PoI through a dynamic fusion block. Furthermore, the features of PoIs derived from the same query box are aggregated together to update the query feature. Our approach prevents information loss caused by view transformation and eliminates the computation-intensive global attention, making the multi-modal 3D object detector more applicable. We conducted extensive experiments on the nuScenes dataset to evaluate our approach. Remarkably, our PoIFusion achieves 74.9% NDS and 73.4% mAP, setting a state-of-the-art record on the multi-modal 3D object detection benchmark. Codes will be made available via https://djiajunustc.github.io/projects/poifusion.
@@ 1 Introduction
Autonomous vehicles are usually equipped with an array of sensors to facilitate safe driving, among which cameras and LiDARs are the most popular. These two sensors are complementary to each other: cameras provide rich textual and color information, while LiDAR sensors supply precise spatial measurements. The effective camera-LiDAR data fusion is widely regarded as a promising direction to achieve high-quality 3D object detection [2, 34, 42, 64], which has attracted a surge of research interest in the community.
Figure 1: A comparison of (a) fusion with unified view, (b) fusion with global attention, and (c) our PoIFusion. “VT” and “PC” indicate view transformation and point clouds, respectively.
The fundamental challenge of camera-LiDAR data fusion arises from the discrepancy of their representation space (i.e., 2D perspective view versus 3D space). To ameliorate this challenge, one common solution is to transform the image and point cloud representations into a unified bird-eye view [42, 34, 22], as depicted in Figure 1(a), or into 3D space [29]. Another recent approach [2, 69], as shown in Figure 1(b), keeps the representation in its original view and abstracts multi-modal features into object queries [6, 40] with the global attention mechanism [56].
However, both of these approaches have inherent issues. In the unified-view approach, the core component, i.e., view transformation, is often based on monocular depth estimation to lift the 2D image into 3D. However, depth estimation is an error-prone task, with errors having a deleterious effect on any downstream tasks (such as object recognition). Moreover, the direct grid-to-grid fusion [34, 42] loses a significant portion of the original representational strengths which comprises modal-specific information [75].
The second query-based approach can avoid feature ambiguity and information loss by keeping the original view of the feature representation. However, the adoption of global attention to integrating multi-modal features incurs high computation and memory overhead. For instance, the state-of-the-art algorithm [69], which follows the query-based approach of fusion with global attention, relies on the well-optimized Flash Attention operator [15] to cut down the time and memory consumption. The high overhead has become an obstacle that hinders the wide application of the algorithm. Furthermore, it is difficult to extract the object-relevant feature with the global attention mechanism [84, 21], especially when it comes to such a large 3D space like the autonomous driving scenario.
@@
As such, we propose a new query-based multi-modal 3D object detection framework that initializes object queries as learnable 3D boxes and dynamically integrates multi-modal features at representative points derived from each object query. In this manuscript, these representative points are referred to as Points of Interest (PoIs), and our framework is named PoIFusion. Intuitively, a naive way to represent a 3D query box with points is to use the center point [8, 69, 40]. However, simply representing a 3D box with its center point totally ignores the geometric properties, such as the size and rotation angle. For supplementary, an improved design is to involve the corners [49]. Nevertheless, sampling multi-modal features according to the projected location of center and corner points also incurs the problem of feature misalignment, since the projection of a 3D box may not be a tight bounding box onto the image view. The issue remains even if a box is well-located in the 3D space [5], not to mention that the query box is not guaranteed to be accurately localized. To ameliorate this problem, our PoIs are adaptively generated from the center and corner points, with box-level and point-level transformation parameters online predicted according to the query feature. In our design, a single PoI serves as the basic unit for fine-grained multi-modal feature fusion, and the ensemble of PoIs derived from the same object query represents the regional feature in a flexible way.
@@
We benchmark the proposed PoIFusion framework on the nuScene dataset and conduct extensive experiments to validate our designs. Notably, our method sets a new record on this challenging dataset, achieving 74.9% NDS and 73.4% mAP on the test set without bells and whistles.
In summary, our contributions are three-fold:
@@
  • We conduct extensive experiments to validate the effectiveness of our method, demonstrating its potential to serve as a strong baseline for future investigation.
@@ 2 Related Work
LiDAR-Based 3D Object Detection.
@@
LiDAR sensors capture 3D point clouds, which provide accurate spatial information that can be important for 3D object detection. Broadly, 3D object detection algorithms operated on point clouds can be categorized into two groups: point-based and voxel-based ones. Point-based 3D object detection algorithms [51, 74, 73] make direct use of the precise coordinates of points, progressively sampling keypoints and extracting local information with set abstraction operators [46, 47, 52, 53] that aggregate information form a point and its neighbours. In contrast, voxel-based 3D object detection algorithms [80, 72, 28, 16, 76, 50, 20, 17] first “voxelize” the point cloud by binning points into a regulare grid. The voxel representation enables straightforward feature extraction using standard (or sparse) convolutions [70, 9, 10] or sparse voxel Transformers [19, 59, 81, 27, 12, 18, 44].
@@
Figure 2: An overview of our proposed PoIFusion framework, which is mainly composed of (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder. In our PoIFusion, we independently extract the feature of each modality and keep the original representation view (i.e., image feature in the perspective view, and point cloud feature in the bird-eye view). The multi-modal feature maps are taken as the input of our multi-modal object decoder. The object decoder is iteratively applied 6 times to integrate the multi-modal feature sampled by generated points of interest (PoIs) and refine the object queries.
Camera-Based 3D Object Detection.
@@
3D object detection is one of the oldest Computer Vision problems and a variety of approaches have been proposed over many years (we do not survey these here). For the purposes of 3D object detection from a moving car a popular approach, and one that can be readily fused with point-cloud estimates, is to measure detection success in the Birds-eye View (BEV) space. Earlier attempts along these lines – e.g., geometric uncertainty [43] and pseudo-LiDAR representation [62] – mainly focus on monocular 3D object detection [61, 48, 3, 14, 79]. Autonomous driving vehicles are typically equipped with multiple cameras, providing perception information over the full 360 degrees. To leverage the relationship between multi-view images, BEV-based and query-based algorithms have been explored. BEV-based 3D object detection [25, 33, 32, 71] explicitly performs view transformation to unify multi-view images into bird-eye-view representation. Query-based 3D object detection [40, 63, 38, 65, 54, 68] follows the pipeline of DETR [6], capitalizing on learnable object queries to extract multi-view information without view transformation.
Multi-Modal 3D Object Detection.
@@
Although the exploration of each individual modality has made encouraging progress, the accuracy and robustness of detection algorithms are still insufficient for safe driving. The fact that images and point clouds are naturally complementary to each other (i.e., rich semantic information versus precise spatial information) has motivated further exploration in multi-modal 3D object detection [60, 2, 8, 22, 67, 31, 26]. In the early stage, the multi-modal approaches utilize point clouds as the principal component, introducing image features at the point level [58, 57, 77, 13] or proposal level [83, 7, 78] to enhance the features of the point clouds. Recently, inspired by multi-view 3D object detection, a series of works [34, 42, 29] propose that unifying the representation space with explicit view transformation [45] facilitates multi-modal fusion. Another group of methods [30, 2, 69, 75] leverages the attention mechanism in Transformer architecture [56] to perform multi-modal fusion in a sequential [2] or a parallel [69] manner. A recent work, ObjectFusion [5], fuses multi-modal features in a two-stage pipeline. It first generates region proposals with the image-augmented BEV features [2], and then extracts the object-centric feature [24, 16] from the voxel, image, and BEV space for further fusion and refinement. In this work, the proposed PoIFusion is also object-centric. However, in contrast to integrating the corresponding region-level feature from multiple modalities, we utilize the PoIs as the basic units to perform multi-modal fusion, which is adaptively derived from the query box - this element stands as the cornerstone of our approach, fundamentally improving the efficacy and flexibility of our framework.
@@ 3.1 Overview
As illustrated in Figure 2, our PoIFusion consists of three main components: (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder. Given multi-view images and point clouds, perspective-view image feature maps $\bm{F}_{I}$ and bird-eye-view (BEV) point cloud feature maps $\bm{F}_{P}$ are independently extracted with the camera encoder and the LiDAR encoder. The camera encoder is composed of an image backbone network and an FPN [35] to obtain multi-scale features. The LiDAR encoder exploits the VoxelNet [70, 80] and a BEV backbone network, following the common practice of the voxel-based paradigm. Then, the multi-modal object decoder, which works as the core component in our method, is iteratively applied 6 times, progressively integrating the multi-modal features and refining the detection boxes.
In our multi-modal decoder, each object query $\bm{Q}$ is formulated as an adaptive 3D box $\bm{b}_{t}$, associating with a feature vector $\bm{v}_{t}$ ($t$ indicates the iteration time step). In each iteration, the object queries are first fed into a self-attention layer to capture the relationships and dependencies between different objects. Here, we follow [38] to exploit a distance-biased self-attention layer in our decoder. Subsequently, for each object query, the box-wise transformation parameters and point-wise shift parameters are generated based on the query feature $\bm{v}_{t}$, and applied on the query box $\bm{b}_{t}$ to obtain a set of PoIs $\bm{P}_{t}=\{P^{i}\}$. To perform multi-modal fusion, a three-step process is undertaken for PoIs. Firstly, each PoI is projected onto both the perspective view and the bird-eye view, establishing the correspondence between the PoI and the multi-modal feature maps. Secondly, the image feature $\bm{f}_{I}$ and point cloud feature $\bm{f}_{P}$ of a PoI are sampled through bilinear interpolation, regarding the projected location on the corresponding view. Thirdly, the extracted multi-modal features at each PoI are dynamically integrated using a dynamic fusion block, followed by a feedforward network (FFN). Finally, a classification head and a regression head are applied to obtain the detection boxes, and the query feature and query box are updated, respectively.
@@ 3.2 PoI Generation
$y_{c}^{\prime}$ $=y_{c}+t_{y},$ $l^{\prime}$ $=l\cdot\exp(t_{l}),$
$z_{c}^{\prime}$ $=z_{c}+t_{z},$ $h^{\prime}$ $=h\cdot\exp(t_{h}),$
$\sin\theta^{\prime}$ $=\sin\theta+t_{\sin},$ $\cos\theta^{\prime}$ $=\cos\theta+t_{\cos}.$
Once the box transformation has been applied, the center point and 8 corner points of $\bm{b^{\prime}}$ are collected together as the anchor points $\{A^{i}\}_{i=0}^{8}$. Finally, the point-wise shift is independently conducted on each anchor point $A^{i}$ to produce a PoI $P^{i}$ ($P^{i}=A^{i}+\Delta_{P}^{i}$). It’s worth noting that one object query corresponds to a set of PoIs derived from the center and corner points of the query box.

### v2

# Feras Dayoub

@@ Abstract
In this work, we present PoIFusion, a conceptually simple yet effective multi-modal 3D object detection framework to fuse the information of RGB images and LiDAR point clouds at the points of interest (PoIs). Different from the most accurate methods to date that transform multi-sensor data into a unified view or leverage the global attention mechanism to facilitate fusion, our approach maintains the view of each modality and obtains multi-modal features by computation-friendly projection and interpolation. In particular, our PoIFusion follows the paradigm of query-based object detection, formulating object queries as dynamic 3D boxes and generating a set of PoIs based on each query box. The PoIs serve as the keypoints to represent a 3D object and play the role of the basic units in multi-modal fusion. Specifically, we project PoIs into the view of each modality to sample the corresponding feature and integrate the multi-modal features at each PoI through a dynamic fusion block. Furthermore, the features of PoIs derived from the same query box are aggregated together to update the query feature. Our approach prevents information loss caused by view transformation and eliminates the computation-intensive global attention, making the multi-modal 3D object detector more applicable. We conducted extensive experiments on nuScenes and Argoverse2 datasets to evaluate our approach. Remarkably, the proposed approach achieves state-of-the-art results on both datasets without any bells and whistles, i.e., 74.9% NDS and 73.4% mAP on nuScenes, and 31.6% CDS and 40.6% mAP on Argoverse2. The code will be made available at https://djiajunustc.github.io/projects/poifusion.
Keywords:
3D Object Detection Autonomous Driving Multi-Sensor Fusion
@@ 1 Introduction
Figure 1: A comparison between the representative multi-modal fusion mechanism in the literature and ours: (a) fusion with a unified view, (b) fusion with global attention, and (c) our proposed fusion at points of interest. “View Trans.”: view transformation. “PoI”: points of interest.
Autonomous vehicles are usually equipped with an array of sensors to facilitate safe driving, among which cameras and LiDARs are the most popular. These two sensors are complementary to each other: cameras provide rich textual and color information, while LiDAR sensors supply precise spatial measurements. The effective camera-LiDAR data fusion is widely regarded as a promising direction to achieve high-quality 3D object detection Bai et al. (2022); Liang et al. (2022); Liu et al. (2023b); Wang et al. (2023d); Zhang et al. (2024); Mao et al. (2023), which has attracted a surge of research interest in the community.
The fundamental challenge of camera-LiDAR data fusion arises from the discrepancy of their representation space (i.e., 2D perspective view versus 3D space). To ameliorate this challenge, one common solution is to transform the image and point cloud representations into a unified bird-eye view Liu et al. (2023b); Liang et al. (2022); Ge et al. (2023), as depicted in Figure 1(a), or into 3D space Li et al. (2022a). Another recent approach Bai et al. (2022); Yan et al. (2023), as shown in Figure 1(b), keeps the representation in its original view and abstracts multi-modal features into object queries Carion et al. (2020); Liu et al. (2022) with the global attention mechanism Vaswani et al. (2017).
However, both of these approaches have inherent issues. In the unified-view approach, the core component, i.e., view transformation, is often based on monocular depth estimation to lift the 2D image into 3D. However, depth estimation is an error-prone task, with errors having a deleterious effect on any downstream tasks (such as object recognition). Moreover, the direct grid-to-grid fusion Liang et al. (2022); Liu et al. (2023b) loses a significant portion of the original representational strengths which comprises modal-specific information Yang et al. (2022c).
The second query-based approach can avoid feature ambiguity and information loss by keeping the original view of the feature representation. However, the adoption of global attention to integrating multi-modal features incurs high computation and memory overhead. For instance, the state-of-the-art algorithm Yan et al. (2023), which follows the query-based approach of fusion with global attention, relies on the well-optimized Flash Attention operator Dao et al. (2022) to cut down the time and memory consumption. The high overhead has become an obstacle that hinders the wide application of the algorithm. Furthermore, it is difficult to extract the object-relevant feature with the global attention mechanism Zhu et al. (2020); Gao et al. (2022), especially when it comes to such a large 3D space like the autonomous driving scenario.
@@
As such, we propose a new query-based multi-modal 3D object detection framework that initializes object queries as learnable 3D boxes and dynamically integrates multi-modal features at representative points derived from each object query. In this manuscript, these representative points are referred to as Points of Interest (PoIs), and our framework is named PoIFusion. Intuitively, a naive way to represent a 3D query box with points is to use the center point Chen et al. (2022a); Yan et al. (2023); Liu et al. (2022). However, simply representing a 3D box with its center point totally ignores the geometric properties, such as the size and rotation angle. For supplementary, an improved design is to involve the corners Sheng et al. (2021). Nevertheless, sampling multi-modal features according to the projected location of center and corner points also incurs the problem of feature misalignment, since the projection of a 3D box may not be a tight bounding box onto the image view. The issue remains even if a box is well-located in the 3D space Cai et al. (2023), not to mention that the query box is not guaranteed to be accurately localized. To ameliorate this problem, our PoIs are adaptively generated from the center and corner points, with box-level and point-level transformation parameters online predicted according to the query feature. In our design, a single PoI serves as the basic unit for fine-grained multi-modal feature fusion, and the ensemble of PoIs derived from the same object query represents the regional feature in a flexible way.
@@
We evaluate the proposed PoIFusion framework on the nuScenes and Argoverse2 datasets, conducting a comprehensive set of experimental analyses to validate our design choices. Notably, PoIFusion achieves state-of-the-art performance on the highly competitive nuScenes benchmark, attaining 74.9% NDS and 73.4% mAP on the test set without any bells and whistles. Moreover, when applied to the more challenging Argoverse2 dataset, our method achieves 40.6% mAP and 31.6% CDS, improving the best performance in the literature by absolutely 4.5% mAP and 3.8% CDS.
In summary, we make three-fold our contributions:
@@
  • We conduct extensive experiments to validate the effectiveness of our method, demonstrating its potential to serve as a strong baseline for this field.
@@ 2 Related Work
2.1 LiDAR-Based 3D Object Detection.
@@
LiDAR sensors capture 3D point clouds, which provide accurate spatial information that can be important for 3D object detection. Broadly, 3D object detection algorithms operated on point clouds can be categorized into two groups: point-based and voxel-based ones. Point-based 3D object detection algorithms Shi et al. (2019); Yang et al. (2020, 2019) make direct use of the precise coordinates of points, progressively sampling keypoints and extracting local information with set abstraction operators Qi et al. (2017a, b); Shi et al. (2020a, 2023) that aggregate information form a point and its neighbors. In contrast, voxel-based 3D object detection algorithms Zhou and Tuzel (2018); Yang et al. (2022a); Lang et al. (2019); Deng et al. (2021a); Yin et al. (2021a); Shi et al. (2022); Fan et al. (2022b); Deng et al. (2021b); Yang et al. (2022b); Wang et al. (2023c, 2024) first “voxelize” the point cloud by binning points into a regular grid. The voxel representation enables straightforward feature extraction using standard (or sparse) convolutions Yan et al. (2018); Chen et al. (2022b, c); Shi et al. (2020b) or sparse voxel Transformers Fan et al. (2022a); Wang et al. (2023a); Zhou et al. (2022); Lai et al. (2023); Chen et al. (2023b); Dong et al. (2022); Mao et al. (2021).
@@
Figure 2: An overview of our proposed PoIFusion framework, which is mainly composed of (a) a camera feature module, (b) a LiDAR feature module, and (c) a PoI multi-modal decoder. In our method, we first independently extract the feature of each modality and keep the original representation view (i.e., image feature in the perspective view, and point cloud feature in the bird-eye view). The multi-modal feature maps are taken as the input of our PoI multi-modal decoder. The decoder is iteratively applied 6 times to integrate the multi-modal feature sampled with generated points of interest (PoIs) and to refine the object queries. In this figure, “proj.” stands for “projection”.
2.2 Camera-Based 3D Object Detection.
@@
3D object detection is one of the oldest Computer Vision problems and a variety of approaches have been proposed over many years (we do not survey these here). For the purposes of 3D object detection from a moving car a popular approach, and one that can be readily fused with point-cloud estimates, is to measure detection success in the Birds-eye View (BEV) space. Earlier attempts along these lines – e.g., geometric uncertainty Lu et al. (2021) and pseudo-LiDAR representation Wang et al. (2019) – mainly focus on monocular 3D object detection Wang et al. (2021b); Reading et al. (2021); Brazil and Liu (2019); Chong et al. (2022); Zhang et al. (2023); Zhou et al. (2021). Autonomous driving vehicles are typically equipped with multiple cameras Xie et al. (2021); Guo et al. (2024), providing perception information over the full 360 degrees. To leverage the relationship between multi-view images, BEV-based and query-based algorithms have been explored. BEV-based 3D object detection Huang et al. (2021); Li et al. (2022c, 2023b, 2023a) explicitly performs view transformation to unify multi-view images into bird-eye-view representation. Query-based 3D object detection Liu et al. (2022); Wang et al. (2022); Liu et al. (2023a); Wang et al. (2023e); Shu et al. (2023); Xiong et al. (2023) follows the pipeline of DETR Carion et al. (2020), capitalizing on object queries to extract multi-view information without view transformation.
2.3 Multi-Modal 3D Object Detection.
@@
Although the exploration of each individual modality has made encouraging progress, the accuracy and robustness of detection algorithms are still insufficient for safe driving. The fact that images and point clouds are naturally complementary to each other (i.e., rich semantic information versus precise spatial information) has motivated further exploration in multi-modal 3D object detection Wang et al. (2023b); Bai et al. (2022); Chen et al. (2022a); Ge et al. (2023); Xie et al. (2023); Li et al. (2024); Jiao et al. (2023); Yin et al. (2024). In the early stage, the multi-modal approaches utilize point clouds as the principal component, introducing image features at the point level Wang et al. (2021a); Vora et al. (2020); Yin et al. (2021b); Chen et al. (2022d) or proposal level Zhu et al. (2022); Chen et al. (2017); Yoo et al. (2020) to enhance the features of the point clouds. Recently, inspired by multi-view 3D object detection, a series of works Liang et al. (2022); Liu et al. (2023b); Li et al. (2022a) propose that unifying the representation space with explicit view transformation Philion and Fidler (2020) facilitates multi-modal fusion. Another group of methods Li et al. (2022b); Bai et al. (2022); Yan et al. (2023); Yang et al. (2022c) leverages the attention mechanism in Transformer architecture Vaswani et al. (2017) to perform multi-modal fusion in a sequential Bai et al. (2022) or a parallel Yan et al. (2023) manner. A recent work, ObjectFusion Cai et al. (2023), fuses multi-modal features in a two-stage pipeline. It first generates region proposals with the image-augmented BEV features Bai et al. (2022), and then extracts the object-centric feature He et al. (2017); Deng et al. (2021a) from the voxel, image, and BEV space for further fusion and refinement.
In this work, the proposed PoIFusion scheme is also object-centric. However, in contrast to integrating the corresponding region-wise feature from multiple modalities, we adaptively generate PoIs from each object query, and leverage the PoIs as the basic units to perform multi-modal fusion - this element stands as the cornerstone of our approach, fundamentally improving the efficacy and flexibility of our multi-modal 3D object detection framework.
@@ 3.1 Overview
As illustrated in Figure 2, our PoIFusion consists of three main components: (a) a camera feature module, (b) a LiDAR feature module, and (c) a PoI multi-modal decoder. Given multi-view images and point clouds, perspective-view image feature maps $\bm{F}_{I}$ and bird-eye-view (BEV) point cloud feature maps $\bm{F}_{P}$ are independently extracted with the image encoder and the voxel encoder. The image encoder is composed of an image backbone network and an FPN Lin et al. (2017a) to obtain multi-scale features. The voxel encoder exploits a sparse 3D backbone network Yan et al. (2018); Zhou and Tuzel (2018) and a BEV backbone network, following the common practice of the voxel-based paradigm. Then, the PoI multi-modal decoder, which works as the core component in our method, is iteratively applied 6 times, progressively integrating the multi-modal features and refining the detection boxes.
In our multi-modal decoder, each object query $\bm{Q}$ is formulated as an adaptive 3D box $\bm{b}_{t}$, associating with a feature vector $\bm{v}_{t}$ ($t$ indicates the iteration time step). In each iteration, the object queries are first fed into a self-attention layer to capture the relationships and dependencies between different objects. Here, we follow Liu et al. (2023a) to exploit a distance-biased self-attention layer in our decoder. Subsequently, for each object query, the box-wise transformation parameters and point-wise shift parameters are generated based on the query feature $\bm{v}_{t}$, and applied on the query box $\bm{b}_{t}$ to obtain a set of PoIs $\bm{P}_{t}=\{P^{i}\}$. To perform multi-modal fusion, a three-step process is undertaken for PoIs. Firstly, each PoI is projected onto both the perspective view and the bird-eye view, establishing the correspondence between the PoI and the multi-modal feature maps. Secondly, the image feature $\bm{f}_{I}$ and point cloud feature $\bm{f}_{P}$ of a PoI are sampled through bilinear interpolation, regarding the projected location on the corresponding view. Thirdly, the extracted multi-modal features at each PoI are dynamically integrated using a dynamic fusion block, followed by a feedforward network (FFN). Finally, a classification head and a regression head are applied for prediction, and the query feature and query box are updated, respectively.
3.2 Object Query Initialization.
We formulate each object query as a learnable dynamic 3D box $\bm{b}\in\mathbb{R}^{8}$ and an attached feature vector $\bm{v}\in\mathbb{R}^{256}$. Specifically, the query box is represented as:
$\bm{b}=[x_{c},y_{c},z_{c},w,l,h,sin\theta,cos\theta],$ (1)
where [$x_{c}$, $y_{c}$, $z_{c}$] is the center location, [$w$, $l$, $h$] is the box dimension, and $\theta$ is the azimuth angle. Notably, although velocity is also predicted in the detection results, it is not factored into the query boxes.
Prior to training, the object query is initialized as follows: the BEV center location [$x_{c}$, $y_{c}$] is set to be uniformly distributed on the BEV space, and the height center $z_{c}$ is set as 0. Besides, the dimension triplet [$w$, $l$, $h$] of each box is initialized as [6, 3, 2], which is about the average size of objects calculated on the dataset. and the azimuth is initialized as 0. The corresponding attached feature vector is randomly initialized.
@@ 3.3 PoI Generation
$y_{c}^{\prime}$ $=y_{c}+t_{y},$ $l^{\prime}$ $=l\cdot\exp(t_{l}),$ (3)
$z_{c}^{\prime}$ $=z_{c}+t_{z},$ $h^{\prime}$ $=h\cdot\exp(t_{h}),$ (4)
$\sin\theta^{\prime}$ $=\sin\theta+t_{\sin},$ $\cos\theta^{\prime}$ $=\cos\theta+t_{\cos}.$ (5)
Once the box transformation has been applied, the center point and 8 corner points of the transformed box $\bm{b^{\prime}}$ are collected together as the anchor points $\{A^{i}\}_{i=0}^{8}$. Finally, the point-wise shift is independently applied to each anchor point $A^{i}$ to produce a PoI $P^{i}$ ($P^{i}=A^{i}+\Delta_{P}^{i}$). It’s worth noting that one object query corresponds to a set of PoIs derived from the center and corner points of the query box.